# centertableaux,boxsize=.5em

### v1

@@ 2.2.1 Example of modified fusion rules from $i^{A}$
Such coset symmetry fusion rule is different from the “untwisted” coset symmetry with $\omega_{D+1}=0$, where the coset symmetry obeys $\mathbb{Z}_{2}$ fusion rule up to tensor product with $2\text{Rep}(\mathbb{Z}_{2}^{2})$.
@@ 2.2.2 Example of modified fusion rules from $i^{B}$
which represents the response of the $\mathbb{Z}_{2}$ Levin-Gu SPT in (2+1)D. In other words, the 1-form symmetry in the bulk generated by the magnetic defect for $a_{1}$ becomes the central extension $\mathbb{Z}_{2}\to\mathbb{Z}_{4}\to\mathbb{Z}_{2}$ extended by the gauged SPT defect. Let us denote the excitation by $s_{\text{LG}}$ stands for the Levin-Gu SPT, then the fusion rule is
@@ 2.3 Effect of $\alpha_{D}$ on coset symmetry
Nontrivial $\alpha_{D}$ for the dynamical $K$ gauge field implies that the $K$-SPT defects given by $i_{k}\alpha_{D}$ are trivial [35].444 Note that in the coset symmetry, the $K$ gauge field does not couple to the rest of the system, only the $G$ defects. Since the $K$-SPT defect is the condensation defect of Wilson lines in the coset symmetry, this can change the coset symmetry fusion rules.
Different choices of $\alpha_{D}$ is related by $\alpha^{\prime}_{D}=\alpha_{D}+\eta_{D}$ with $\eta_{D}\in H^{D}(BK,U(1))$ the $K$ SPT response. When $\eta_{D}$ is obtained by restricting $G$ SPT response to the subgroup $K\subset G$, this correspond to fusing the boundary with a gauged SPT defect for $G$ in the bulk. In that case, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ does not modify the symmetry category, as the gauged SPT defect induces the automorphism of the topological operators at the boundary. However, note that the shift by $\eta_{D}$ in general do not have such bulk interpretation. When $\eta_{D}$ is not obtained by the bulk gauged SPT defect, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ can modify the symmetry category.
The boundary action $\alpha_{D}$ means that when a magnetic defect for $K$ terminates on the boundary to give rise to dynamical $K$ gauge field holonomies on the boundary, the ending locus has additional gauged SPT defects for the $K$ remnant gauge group as discussed in [35]. In other words, the symmetry generated by these gauged SPT defects for $K$ can terminate on the condensed magnetic flux, and different choices of $\alpha_{D}$ implies that different subgroup symmetry of the gauged SPT defects for $K$ can terminate and become trivial. Such change of condensed objects by shifting $\alpha\to\alpha_{D}+\eta_{D}$ modifies the symmetry category, as we will see below with an example.
@@ 2.3.1 Example of modified fusion rules from $\alpha_{D}$
The SPT action $\alpha_{2}$ shifts the condensed particles; it exchanges the particles by $m^{2}\to m^{2}e^{\prime},m^{\prime}\to m^{\prime}e$ at the boundary, so the condensed particles are now generated by $m^{2}e^{\prime},e^{2},m^{\prime}e$. Note that this action of $\alpha_{2}$ on condensed anyons is not an automorphism of the bulk $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ gauge theory, and $\alpha_{2}$ cannot be obtained from the gauged SPT defect in the bulk. The symmetry at the boundary is still generated by the lines of $m,e,e^{\prime}$, but they now generate the non-anomalous $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ symmetry, as $m^{2}\sim e^{\prime}$ up to condensed particles. This shows that different choices of $\alpha_{2}$ can modify the fusion rule of the coset symmetry.
@@ 2.4 Modified Frobenius-Schur indicator from twist
The generalized Frobenius-Schur indicator or the associator for an invertible $G$ symmetry in $D$ spacetime dimensions corresponds to $(D+1)$-cocycle $\omega_{D+1}$ for group $G$, and they are phase factors that relate different ways of fusing $(D+1)$ symmetry defects for $G$. Such a cocycle corresponds to an anomaly for the $G$ symmetry (both an obstruction to realization in SPT phases and obstruction to gauging the $G$ symmetry). The bulk TQFT description for the Frobenius-Schur indicator is the Boltzmann weight given by the phase factor associated with intersection of $(D+1)$ symmetry defects at a point.
@@ 2.6.2 Critical points with twisted coset symmetry
Consider a 2+1D massless free fermion in the fundamental representation of a finite group $G$ symmetry. Due to the parity anomaly, the theory is a gapless “anomalous SPT” phase with $G$ symmetry. Let us then gauge a non-normal subgroup $K\subset G$, and we choose $G,K$ such that the anomalous SPT phase restricted to $K$ is trivial. This gives a gapless system with twisted coset symmetry $(G,K,\omega_{D+1},\alpha_{D})$, where $\omega_{D+1}$ is exact for the entire $G$ (since it is an anomalous SPT).
As an example, consider a free fermion theory with $S_{3}=\mathbb{Z}_{3}\rtimes\mathbb{Z}_{2}$ symmetry. For example, it can be two Dirac fermions with charge ±$\pm 1$ under $\mathbb{Z}_{3}$ symmetry, and there is $\mathbb{Z}_{2}$ symmetry that exchange them and also complex conjugate the fermions. In terms of the Majorana components $\psi_{11}+i\psi_{12},\psi_{21}+i\psi_{22}$, the $\mathbb{Z}_{2}$ symmetry acts as $\psi_{11}\leftrightarrow\psi_{21}$, $\psi_{12}\leftrightarrow-\psi_{22}$, i.e. a doublet complex fermion of $\mathbb{Z}_{2}$ given by $(\psi_{11}-\psi_{21})+i(\psi_{12}+\psi_{22})$ and a singlet complex fermion.
We will consider $\mathbb{Z}_{3},\mathbb{Z}_{2}$ symmetry independently, since $H^{3}(BS_{3},U(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{3}$ factorizes into the $\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$ parts. Let us use the regularization such that fermion with positive masses describe trivial SPT phases.
The system enriched with $\mathbb{Z}_{3}$ symmetry with negative masses of the fermions flow to the SPT phase
$-\frac{2}{4\pi}AdA-4\text{CS}_{\text{grav}}~{},$ (2.23)
where $A$ is the $\mathbb{Z}_{3}$ gauge field with holonomy in $\frac{2\pi}{3}\mathbb{Z}$. In the following, we will omit the gravitational Chern-Simons term $\text{CS}_{\text{grav}}$.555 More explicitly, $\int\text{CS}_{\text{grav}}=\frac{1}{192\pi}\int\text{Tr}\left(\omega d\omega+\frac{2}{3}\omega^{3}\right)~{},$ (2.24) where $\omega$ is the spin connection. The SPT phase corresponds to $\mathbb{Z}_{3}$ generator in $H^{3}(BS_{3},U(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{3}$.
The system enriched with $\mathbb{Z}_{2}$ symmetry with negative masses of the fermions flow to nontrivial SPT phase due to the single complex fermion in the sign representation of $\mathbb{Z}_{2}$:
$-\frac{1}{4\pi}A^{\prime}dA^{\prime}~{},$ (2.25)
where $A^{\prime}$ is the $\mathbb{Z}_{2}$ gauge field with holonomy in $\pi\mathbb{Z}$, and we have omitted the gravitational Chern-Simons term. The SPT phase corresponds to $\mathbb{Z}_{2}$ generator in $H^{3}(BS_{3},U(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{3}$.
Therefore we conclude that the SPT phase for negative masses is the sum of generator in $H^{3}(BS_{3},U(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{3}$. For massless fermions, the SPT phase is half of the sum, and this is an anomalous SPT phase. In other words, the massless fermions enriched with $S_{3}$ symmetry is a gapless (anomalous) SPT. As a consequence, if we further gauge the $\mathbb{Z}_{2}$ symmetry, this gives twisted coset symmetry $S_{3}/\mathbb{Z}_{2}$:
$\left(G=S_{3},\,K=\mathbb{Z}_{2},\,\omega_{4}=d\eta_{3}/2,\,\alpha_{3}=\frac{1}{2}\eta_{3}\right)~{},$ (2.26)
where $\eta_{3}$ is the $\mathbb{Z}_{2}$ generator in $H^{3}(BS_{3},U(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{3}$.
Since this is a gapless system with twisted coset symmetry, the system is also a gapless SPT with coset symmetry. [^6] For other examples of gapless SPT phases, see e.g.  [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51].
We note that since $\alpha_{D}$ is not properly quantized, this twisted coset symmetry cannot be realized in SPT phases.
@@ 3.1 Anomaly conditions
  • A subgroup $K^{\prime}\subset G$ such that $[\omega_{D+1}|_{K^{\prime}}]=0$.
@@
    The subgroups $K,K^{\prime}$ for the gapped boundaries need to satisfy $[\omega_{D+1}|_{K}]=1,[\omega_{D+1}|_{K^{\prime}}]=1$, i.e., they become exact cocycles under pullback using the inclusion map $K,K^{\prime}\subset G$. This also guarantees that the magnetic operators of $K,K^{\prime}$ holonomy are not attached to gauged SPT operators [35].
@@ 3.2 SPT phases with twisted coset symmetry $(H\Join K)/K$
  • $G$ is expressed as the bicrossed product with some subgroup $H\subset G$ as $G=H\Join K$. Namely, for two subgroups $H,K$ of $G$, $G$ can be expressed as $G=HK$ and $H\cap K=\{\text{id}\}$. Equivalently, each group element $g\in G$ has a unique expression as $g=hk$ with $h\in H,k\in K$ [^7] When $H$ is a normal subgroup, the bicrossed product reduces to the semidirect product $H\rtimes K$.
  • There exists a choice of the above $H\subset G$ satisfying $[\omega_{D+1}|_{H}]=0$.
To check this, let us consider a bulk symmetry TQFT in $(D+1)$D on an interval, where the $G$ gauge theory with the twist $\omega_{D+1}$ is sandwiched by a pair of gapped boundary conditions. One of the gapped boundary breaks the gauge symmetry $G\to K$, while the other breaks as $G\to H$. We note that this SSB pattern $G\to H$ is made possible thanks to the above two conditions. The coset symmetry defects are realized as the topological operators at the boundary $G\to K$. One can see that this system realizes the SPT after shrinking the interval by the following:
@@
Since the gapped boundary $G\to H$ can be additionally twisted by a group cohomology $\eta_{D}\in H^{D}(BH,U(1))$, a subclass of SPT with $(H\Join K)/K$ symmetry in $D$ spacetime dimensions is classified by the group cohomology $H^{D}(BH,U(1))$. For untwisted coset symmetries with $\omega_{D+1}=0$, the SPT can be obtained by starting with the SSB phase with $G\to H$ with the SPT action $\eta_{D}$ for the unbroken symmetry, and then gauging the broken $K$ symmetry. An explicit lattice model for an SPT with untwisted $(H\Join K)/K$ symmetry will be presented in Sec. 4.1.
@@ 3.3 Anomalies of twisted coset symmetry: anomaly from bulk cocycle
The twist $\omega_{D+1}$ in the coset symmetry typically leads to the nontrivial anomalies. To illustrate the obstruction to SPT phase with coset symmetry due to nontrivial twist $\omega_{D+1}$, let us consider the example $G=G_{0}\times G_{1}$ and $K$ is given by a non-normal subgroup of $G_{0}$. Moreover, there is nontrivial $\omega_{D+1}$ given by topological term for $G_{1}$. Then on any gapped boundary, $G_{1}$ needs to break to subgroup of $K_{1}$ such that $\omega_{D+1}|_{K_{1}}=d\alpha$, and in particular the reference boundary $G\to K$ breaks $G_{1}$ completely. This means that there are Wilson lines of $G_{1}$ that can connect the two boundaries and thus violating condition (3) presented in Sec. 3.1. This gives an example of anomalous twisted coset non-invertible symmetry.

### v3

@@ 2.2.1 Example of modified fusion rules from $i^{A}$
Such coset symmetry fusion rule is different from the “untwisted” coset symmetry with $\omega_{D+1}=0$, where the coset symmetry obeys $\mathbb{Z}_{2}$ fusion rule up to tensor product with $\text{Rep}(\mathbb{Z}_{2}^{2})$.
@@ 2.2.2 Example of modified fusion rules from $i^{B}$
which represents the response of the $\mathbb{Z}_{2}$ Levin-Gu SPT in (2+1)D. In other words, the 1-form symmetry in the bulk generated by the magnetic defect for $a_{2}$ becomes the central extension $\mathbb{Z}_{2}\to\mathbb{Z}_{4}\to\mathbb{Z}_{2}$ extended by the gauged SPT defect. Let us denote the excitation by $s_{\text{LG}}$ stands for the Levin-Gu SPT, then the fusion rule is
@@ 2.3 Effect of $\alpha_{D}$ on coset symmetry
Different choices of $\alpha_{D}$ are related by $\alpha^{\prime}_{D}=\alpha_{D}+\eta_{D}$ with $\eta_{D}\in H^{D}(BK,U(1))$ the $K$ SPT response. When $\eta_{D}$ is obtained by restricting $G$ SPT response to the subgroup $K\subset G$, this correspond to fusing the boundary with a gauged SPT defect for $G$ in the bulk. In that case, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ does not modify the symmetry category, as the gauged SPT defect induces the automorphism of the topological operators at the boundary. However, note that the shift by $\eta_{D}$ in general does not have such bulk interpretation. When $\eta_{D}$ is not obtained by the bulk gauged SPT defect, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ can modify the symmetry category.
The boundary action $\alpha_{D}$ means that when a magnetic defect for $K$ terminates on the boundary to give rise to dynamical $K$ gauge field holonomies on the boundary, the ending locus has additional gauged SPT defects for the $K$ remnant gauge group as discussed in [38]. Such changes of condensed objects by shifting $\alpha\to\alpha_{D}+\eta_{D}$ modify the symmetry category.
Concretely, let us consider the group of center magnetic defects in the bulk $Z_{\omega}(G)$, as defined in (2.11). The bulk center symmetry $Z_{\omega}(G)$ is given by the central extension of the coset symmetry $Z_{\omega}(G/K)$ at the boundary,
$Z_{\omega}(G)\cap K\to Z_{\omega}(G)\to Z_{\omega}(G/K)~.$ (2.21)
Let us write the above extension class $\nu\in H^{2}(Z_{\omega}(G/K),Z_{\omega}(G)\cap K)$. While in Sec. 2.2.2 we have seen that the coset symmetry $Z_{\omega}(G/K)$ gets extended due to the bulk cocycle $\omega_{D+1}$ through the slant product $i^{B}$, the boundary cochain $\alpha_{D}$ also extends the coset symmetry through its slant product. This is because the gauged $K$ SPT defects generally get identified with the $K$ magnetic defects at boundary up to condensed objects, meaning that the gauged $K$ SPT defects replace the group $Z_{\omega}(G)\cap K$ in (2.21) in the structure of coset symmetry.
The coset symmetry has the form of the central extension with the gauged $K$ SPT defect,
$1\rightarrow H^{D-1}(BK,U(1))\rightarrow{\cal G}^{(0)}_{G/K}\rightarrow Z_{\omega}(G/K)\rightarrow 1~,$ (2.22)
whose extension class is characterized by
$\Omega(\tilde{g},\tilde{g}^{\prime})=\frac{1}{|\omega_{D+1}|}\left(i_{(\tilde{g},0)}^{B}\omega_{D+1}+i_{(\tilde{g}^{\prime},0)}^{B}\omega_{D+1}-i_{(\tilde{g}+\tilde{g}^{\prime},\nu(\tilde{g},\tilde{g}^{\prime}))}^{B}\omega_{D+1}\right)|_{K}+i_{\nu(\tilde{g},\tilde{g}^{\prime})}\alpha_{D}~,$ (2.23)
where $\tilde{g},\tilde{g}^{\prime}\in Z_{\omega}(G/K)$, and we label the elements of $Z_{\omega}(G)$ by a pair $(\tilde{g},\tilde{k})\in Z_{\omega}(G/K)\times(Z_{\omega}(G)\cap K)$. Also $\mid_{K}$ means the restriction to the group $K$. Below, we will provide an example where $\alpha_{D}$ modifies the fusion rule by the extension.
@@ 2.3.1 Example of modified fusion rules from $\alpha_{D}$
The SPT action $\alpha_{2}$ shifts the condensed particles; it exchanges the particles by $m^{2}\to m^{2}e^{\prime},m^{\prime}\to m^{\prime}e$ at the boundary, so the condensed particles are now generated by $m^{2}e^{\prime},e^{2},m^{\prime}e$. Note that this action of $\alpha_{2}$ on condensed anyons is not an automorphism of the bulk $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ gauge theory, and $\alpha_{2}$ cannot be obtained from the gauged SPT defect in the bulk. The symmetry at the boundary is still generated by the lines of $m,e,e^{\prime}$, but they now generate the non-anomalous $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ symmetry, as $m^{2}\sim e^{\prime}$ up to condensed particles. We note that the electric particle $e^{\prime}$ gives the central extension of the coset symmetry $m$ from $\mathbb{Z}_{2}$ to $\mathbb{Z}_{4}$; $e^{\prime}$ corresponds to the slant product of $\alpha_{2}$ shown in the description of (2.23). This shows that different choices of $\alpha_{2}$ can modify the fusion rule of the coset symmetry.
@@ 2.4 Modified Frobenius-Schur indicator from twist
As discussed in [39], the Frobenius-Schur indicator $\kappa_{a}$ is a $\mathbb{Z}_{2}$ valued piece of data associated with self-dual topological lines $a=\bar{a}$. It is part of the $F$ symbols for such topological lines, in particular $F^{a\bar{a}a}_{a}]_{11}=[F^{aaa}_{a}]_{11}$ where $1$ is the vacuum object. There is also a Frobenius-Schur indicator as part of the associator data for 0-form symmetries in 1+1d. These two Frobenius-Schur indicators are related to each other: the topological defects whose fusion are modified by the Frobenius-Schur indicator in 1+1d can be thought of as living on the boundary of a $\mathbb{Z}_{2}$ 0-form symmetry domain wall that extends to 2+1d bulk, and the Frobenius-Schur indicator is given by whether the $\mathbb{Z}_{2}$ symmetry has nontrivial bosonic SPT (see e.g. [6, 7]). After gauging the $\mathbb{Z}_{2}$ symmetry in the bulk, the $\mathbb{Z}_{2}$ symmetry fluxes give rise to new self-dual topological lines, whose associators carry information about the bosonic $\mathbb{Z}_{2}$ SPT. Such SPTs can be described a 3-cocycle for group $\mathbb{Z}_{2}$ [40]. There is only one nontrivial such cocycle, corresponding to the Levin-Gu phase [41], and this is consistent with the Frobenius-Schur indicator being $\mathbb{Z}_{2}$ valued in 2+1d.
In analogy, in this work, we define a generalized Frobenius-Schur indicator as a particular part of the data describing the associator of topological defects in 3+1d, in 1-1 correspondence to SPTs of the bulk symTFT. For bulk symmetry $G$, the generalized Frobenius-Schur indicator is labeled by $G$-SPT in the bulk, as described by a $(D+1)$-cocycle $\omega_{D+1}$ for group $G$ [40], and they are phase factors that relate different ways of fusing $(D+1)$ symmetry defects for $G$. Such a cocycle corresponds to an anomaly for the $G$ symmetry (both an obstruction to realization in SPT phases and obstruction to gauging the $G$ symmetry). The bulk TQFT description for the Frobenius-Schur indicator is the Boltzmann weight given by the phase factor associated with intersection of $(D+1)$ symmetry defects at a point.
@@ 2.6.2 Critical points with twisted coset symmetry
We will present a critical theory that has fractional response for coset non-invertible symmetry. The example is given by massless Dirac fermion in 2+1d, coupled to $\mathbb{Z}_{2}$ gauge field where the $\mathbb{Z}_{2}$ gauge transformation complex conjugates the Dirac fermion. Before gauging the $\mathbb{Z}_{2}$ symmetry, the theory has $O(2)$ symmetry, and the $O(2)$ symmetry has fractional response
$O(2)_{1/2,1/2}+\text{Dirac fermion}~,$ (2.29)
where the subscript denotes fractional Chern-Simons level for both the continuous and the $\mathbb{Z}_{2}$ part, following the notation in [35]. Since the $\mathbb{Z}_{2}$ symmetry complex conjugates the Dirac fermion that consists of two Majorana fermions, it leaves one of the Majorana fermion invariant and flips the sign of the other Majorana fermion. After gauging the $\mathbb{Z}_{2}$ symmetry, the symmetry is the coset symmetry $O(2)/\mathbb{Z}_{2}$. The coset symmetry has $\alpha=\frac{1}{2}\eta$ where $\eta$ is the effective action for the minimal $\mathbb{Z}_{2}$ Chern-Simons term given by gauging the $\mathbb{Z}_{2}$ symmetry in the root $\mathbb{Z}_{2}$ fermionic SPT phase. The coset symmetry has $\omega$ given by an exact cocycle, which contains fractional response of $O(2)/\mathbb{Z}_{2}$ symmetry.
We can also consider discrete symmetry. For example, take a Dirac fermion, and focus on the $\mathbb{D}_{4n}=\mathbb{Z}_{2n}\rtimes\mathbb{Z}_{2}\subset O(2)=U(1)\rtimes\mathbb{Z}_{2}$ subgroup of the previous example. The previous discussion carries over to give massless fermion with fractional response for the coset $(\mathbb{Z}_{2n}\rtimes\mathbb{Z}_{2})/\mathbb{Z}_{2}$ symmetry.
@@ 3.1 Anomaly conditions
  • A subgroup $K^{\prime}\subset G$ such that $\omega_{D+1}|_{K^{\prime}}]=0$.
@@
    The subgroups $K,K^{\prime}$ for the gapped boundaries need to satisfy $\omega_{D+1}|_{K}]=0,[\omega_{D+1}|_{K^{\prime}}]=0$, i.e., they become exact cocycles under pullback using the inclusion map $K,K^{\prime}\subset G$. This also guarantees that the magnetic operators of $K,K^{\prime}$ holonomy are not attached to gauged SPT operators [38].
@@ 3.2 SPT phases with twisted coset symmetry $(H\Join K)/K$
  • $G$ is expressed as the bicrossed product with some subgroup $H\subset G$ as $G=H\Join K$. Namely, for two subgroups $H,K$ of $G$, $G$ can be expressed as $G=HK$ and $H\cap K=\{\text{id}\}$. Equivalently, each group element $g\in G$ has a unique expression as $g=hk$ with $h\in H,k\in K$. [^4] When $H$ is a normal subgroup, the bicrossed product reduces to the semidirect product $H\rtimes K$.
  • There exists a choice of the above $H\subset G$ satisfying $\omega_{D+1}|_{H}]=0$.
To check this, let us consider a bulk symmetry TQFT in (D+1)(D+1)D on an interval, where the $G$ gauge theory with the twist $\omega_{D+1}$ is sandwiched by a pair of gapped boundary conditions. One of the gapped boundary breaks the gauge symmetry $G\to K$, while the other breaks as $G\to H$. We note that this SSB pattern $G\to H$ is made possible thanks to the above two conditions. The coset symmetry defects are realized as the topological operators at the boundary $G\to K$. One can see that this system realizes the SPT after shrinking the interval by the following:
@@
Since the gapped boundary $G\to H$ can be additionally twisted by a group cohomology $\eta_{D}]\in H^{D}(BH,U(1))$, a subclass of SPT with $(H\Join K)/K$ symmetry in $D$ spacetime dimensions is classified by the group cohomology $H^{D}(BH,U(1))$. For untwisted coset symmetries with $\omega_{D+1}=0$, the SPT can be obtained by starting with the SSB phase with $G\to H$ with the SPT action $\eta_{D}$ for the unbroken symmetry, and then gauging the broken $K$ symmetry. An explicit lattice model for an SPT with untwisted $(H\Join K)/K$ symmetry will be presented in Sec. 4.1.
@@ 3.3 Anomalies of twisted coset symmetry: anomaly from bulk cocycle
The twist $\omega_{D+1}$ in the coset symmetry typically leads to the nontrivial anomalies. To illustrate the obstruction to SPT phase with coset symmetry due to nontrivial twist $\omega_{D+1}$, let us consider the example $G=G_{0}\times G_{1}$ and $K$ is given by a non-normal subgroup of $G_{0}$. Moreover, there is nontrivial $\omega_{D+1}$ given by topological term for $G_{1}$. Then on any gapped boundary, $G_{1}$ needs to break to subgroup $K_{1}$ such that $\omega_{D+1}|_{K_{1}}=d\alpha$, and in particular the reference boundary $G\to K$ breaks $G_{1}$ completely. This means that there are Wilson lines of $G_{1}$ that can connect the two boundaries and thus violating condition (3) presented in Sec. 3.1. This gives an example of anomalous twisted coset non-invertible symmetry.